OpenStudy (anonymous):

Find the distance between the points (2, 6) and (-3, 4). a. √27 units b. √31 units c. √29 units d. √30 units

hero (hero):

c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

distance equals sqrt(a^2 - b^2), where a is the difference of x2-x1 and b is the difference of y2-y1... so a = (-3-2) = -5, b = (4-6) = -2 then sqrt( (-5)^2 + (-2)^2) = sqrt (25+4) or... sqrt(29)